OpenStudy (anonymous):

if RS=5x-7, ST= 2x=8 and RT= 2x-5. what is theperimeter of triangle RST? is it 51, 45, 57, or 47?

OpenStudy (dtan5457):

Is there a given value for x? Seems like there is missing information to solve this

OpenStudy (anonymous):

there is no shown value for x. i got 9x+6 i just dont know how to reach the correct answer of 51.

OpenStudy (anonymous):

angles r & t are congruent

Nnesha (nnesha):

ST = 2x ( plus or minus ) 8 ?

OpenStudy (anonymous):

plus 8

Nnesha (nnesha):

how did you get 9x +6 ?? :o

OpenStudy (anonymous):

by adding the segments together

Nnesha (nnesha):

\[\huge\rm 5x-7+2x +8 +2x -5 \cancel=9x +6\]

Nnesha (nnesha):

-7+8-5 =?

OpenStudy (nikato):



OpenStudy (anonymous):

-4?

Nnesha (nnesha):

yes

OpenStudy (nikato):

since <R=<T, its an isoscles triangle with side RS=ST. so 5x-7=2x+8 and you can solve for x. then just plug this value to whatever expression you get when you add all three sides

OpenStudy (anonymous):

how do i reach 51?

OpenStudy (nikato):

well. now you know the perimeter is 9x-4 and using that equation i gave you above, solve for x

OpenStudy (anonymous):

oh okay i got it, thank you so much!

OpenStudy (nikato):

no problem!